a ball was thrown from height of 2ft at a velocity of 35m/s at 25degrees

a. how long was the ball in the air
b. what is the distance the ball was thrown
c. what is the velocity of the ball as it hits the floor
Answer #1

Vo =35m/s 0 = 250 Por =h Tou ha 2 eft = 0.305 m y = voyt + t ayt? = - h= vosindt-gth -) - 0.305= 35 sinnsot-gift? 4.9t² - 14.